An electric dipole, consisting of charges of magnitude 10.0 nC separated by 15.1 ?m, is in an electric field of strength 1520.0 N/C. What is the magnitude of the electric dipole moment?

What is the difference in potential energy corresponding to dipole orientations parallel to and 45 degrees to the field?
Answer #1


P = q*l = 10e-9*15.1e-6 = 1.51e-13 Cm

---------------

dU = P*E*(1-cos45) = 6.72*10^-11 J
Answer #2

The electric dipole moment is

p = q ( L) = 10 * 10 ^-9 C (15.1 * 10 ^-6 m) = 1.51 * 10 ^ -13 Cm

the difference in potential energy is

del U = Ui- Uf= pE cos theta - pE cos theta = pE ( cos 0-cos 45 )

          =1.51 * 10 ^ -13 Cm(1520 N/C) ( 1- 0.707 )

            =672.4936* 10 ^ -13 J